# The Limits of Assumption-free Tests for Algorithm Performance

### v1

@@ Abstract
Algorithm evaluation and comparison are fundamental questions in machine learning and statistics—how well does an algorithm perform at a given modeling task, and which algorithm performs best? Many methods have been developed to assess algorithm performance, often based around cross-validation type strategies, retraining the algorithm of interest on different subsets of the data and assessing its performance on the held-out data points. Despite the broad use of such procedures, the theoretical properties of these methods are not yet fully understood. In this work, we explore some fundamental limits for answering these questions with limited amounts of data. In particular, we make a distinction between two questions: how good is an algorithm $\mathcal{A}$ at the problem of learning from a training set of size $n$, versus, how good is a particular fitted model produced by running $\mathcal{A}$ on a particular training data set of size $n$?
Our main results prove that, for any test that treats the algorithm $\mathcal{A}$ as a “black box” (i.e., we can only study the behavior of $\mathcal{A}$ empirically), there is a fundamental limit on our ability to carry out inference on the performance of $\mathcal{A}$, unless the number of available data points $N$ is many times larger than the sample size $n$ of interest. (On the other hand, evaluating the performance of a particular fitted model is easy as long as a holdout data set is available—that is, as long as $N-n$ is not too small.) We also ask whether an assumption of algorithmic stability might be sufficient to circumvent this hardness result. Surprisingly, we find that this is not the case: the same hardness result still holds for the problem of evaluating the performance of $\mathcal{A}$, aside from a high-stability regime where fitted models are essentially nonrandom. Finally, we also establish similar hardness results for the problem of comparing multiple algorithms.
@@ 1 Introduction
Evaluating the performance of a regression algorithm, and comparing performance of different algorithms, are fundamental questions in machine learning and statistics (Salzberg,, 1997; Dietterich,, 1998; Bouckaert,, 2003; Demsar,, 2006; Garcia and Herrera,, 2008; Hastie et al.,, 2009; Raschka,, 2018). When performing a data analysis, if the properties of the underlying data distribution are unknown, how can we determine which algorithms would model the data well? For instance, for the task of predicting a real-valued response $Y$, which algorithm would have the smallest possible error?
Algorithm evaluation.
@@
EvaluateAlg: how well does algorithm $\mathcal{A}$ perform on data drawn from $P$?
Writing ${\widehat{f}}$ to denote the fitted model produced by running algorithm $\mathcal{A}$ on a training data set, a related question is to ask about the performance of the trained model ${\widehat{f}}$:
EvaluateModel: how well does the fitted model ${\widehat{f}}$ perform on data drawn from $P$?
@@
In practice, these two questions are often treated as interchangeable: for instance, a paper might state that “Logistic regression performs well on this problem” (which suggests that the question EvaluateAlg is being addressed), but then justify the statement by using a holdout set to evaluate the fitted model ${\widehat{f}}$ produced by running logistic regression on a particular data set (which is actually addressing the question EvaluateModel). In this work, we argue that these two questions are not equivalent, and indeed, answering EvaluateAlg may be impossible even in settings where EvaluateModel is trivial to answer.
Algorithm comparison.
@@
CompareAlg: which algorithm, $\mathcal{A}_{0}$ or $\mathcal{A}_{1}$, performs better on data drawn from $P$?
@@
CompareModel: which fitted model, ${\widehat{f}}_{0}$ or ${\widehat{f}}_{1}$, performs better on data drawn from $P$?
@@ 1.1 Problem formulation: evaluating an algorithm
    with $\mathcal{Y}=\widehat{\mathcal{Y}}=\mathbb{R}$.
@@
which is computed with respect to a test point sampled as $(X,Y)\sim P$. Now consider a regression algorithm $\mathcal{A}$, which inputs a data set $\{(X_{i},Y_{i})\}_{i\in[n]}$ and returns a fitted model ${\widehat{f}}_{n}$. The fitted model ${\widehat{f}}_{n}$ has risk $R_{P}({\widehat{f}}_{n})$—this captures the expected loss for the future performance of ${\widehat{f}}_{n}$, on test data drawn from $P$. Note we can write
@@
where now the expected value is taken with respect to a random draw $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X,Y)\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$—that is, with respect to a random draw of both the training and test data. In particular, these two measures of risk are related as
@@
With this notation in place, we can now make our questions, EvaluateAlg and EvaluateModel, more precise.
EvaluateAlg: what is $R_{P,n}(\mathcal{A})$?  versus  EvaluateModel: what is $R_{P}({\widehat{f}}_{n})$?
These two questions are different: a low value of $R_{P,n}(\mathcal{A})$ tells us that $\mathcal{A}$ tends to perform well for distribution $P$, while a low value of $R_{P}({\widehat{f}}_{n})$ tells us that this particular fitted model produced by $\mathcal{A}$ performs well. Which question is the “right” question to ask, will depend on the context, and we do not claim that one of these questions is more important than the other; instead, the aim of our paper is to understand the distinction between these two questions, and to emphasize that the two questions should not be treated interchangeably.
Defining a hypothesis test.
There are many possible goals we might define for addressing EvaluateAlg: for example, we might wish to construct a confidence interval for the risk $R_{P,n}(\mathcal{A})$, or to perform a hypothesis test assessing whether $R_{P,n}(\mathcal{A})$ lies above or below some prespecified threshold $\tau$,
$H_{0}:\,R_{P,n}(\mathcal{A})\geq\tau\textnormal{ \ versus \ }H_{1}:\,R_{P,n}(\mathcal{A})<\tau.$ (1)
Indeed, to establish a hardness result for these types of questions, it suffices to consider the problem of performing the hypothesis test (1). This is because, if we have instead constructed a confidence interval $\widehat{C}$ with a guarantee of containing the risk $R_{P,n}(\mathcal{A})$ with probability $\geq 1-\alpha$, we can test (1) at Type I error level $\alpha$ by simply checking whether $\widehat{C}\subseteq[0,\tau)$.
@@ 1.2 Overview of contributions
In this paper, for the problem of algorithm evaluation, we will establish fundamental limits on the ability of any black-box test (that is, any test that studies the behavior of $\mathcal{A}$ empirically rather than theoretically) to perform inference on the hypothesis test (1). Surprisingly, we will also see that this hardness result is not alleviated by adding an assumption of algorithmic stability. We will also establish analogous results for the problem of algorithm comparison.
Organization.
In Section 2, we will present some background on algorithm testing and on algorithmic stability, in addition to giving an overview of some related work in the literature. Section 3 presents our main results establishing the hardness of testing EvaluateAlg, and establishes that an extremely simple and naive test already achieves (nearly) the maximum possible power. Results under an algorithmic stability assumption are presented in Section 4. We then turn to the problem of algorithm comparison, which is addressed in Section 5. Discussion and conclusions are provided in Section 6. All proofs are deferred to the Appendix.
@@ 2 Background
In this section, we formalize some definitions and give additional background for studying properties of algorithms. Formally, for data lying in some space $\mathcal{X}\times\mathcal{Y}$, we define an algorithm $\mathcal{A}$ as a map
@@
That is, $\mathcal{A}$ takes as input $\{(X_{i},Y_{i})\}_{i\in[n]}$, a data set of any size, and returns a fitted model ${\widehat{f}}_{n}:\mathcal{X}\rightarrow\widehat{\mathcal{Y}}$, which maps a feature value $x$ to a fitted value, or prediction, ${\widehat{y}}={\widehat{f}}_{n}(x)$. For instance, for a real-valued response $Y\in\mathcal{Y}=\mathbb{R}$, we might take $\mathcal{A}$ to be the least squares regression algorithm, leading to linear fitted models ${\widehat{f}}_{n}$. In many settings, commonly used algorithms $\mathcal{A}$ will include some form of randomization—for instance, stochastic gradient descent type methods. To accommodate this setting, we will expand the definition of an algorithm to take an additional argument,
@@
The fitted model ${\widehat{f}}_{n}=\mathcal{A}\big{(}\{(X_{i},Y_{i})\}_{i\in[n]};\xi\big{)}$ is now obtained by running algorithm $\mathcal{A}$ on data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, with randomization provided by the argument $\xi\in[0,1]$, which acts as a random seed. [^1] Throughout this paper, we implicitly assume measurability for all functions we define, as appropriate. In the case of the algorithm $\mathcal{A}$, this takes the following form: we assume measurability of the map $\left((x_{1},y_{1}),\ldots,(x_{n},y_{n}),\xi,x\right)\mapsto[\mathcal{A}(\{(x_{i},y_{i})\}_{i\in[n]};\xi)](x)$. Of course, the general definition (2) of a randomized algorithm also accommodates the deterministic (i.e., non-random) setting—we can simply choose an algorithm $\mathcal{A}$ whose output depends only on the data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, and ignores the second argument $\xi$.
@@ 2.1 Black-box tests
To understand the difficulty of the question EvaluateAlg, we need to formalize the limitations on what information is available to us, as the analyst, for performing inference on an algorithm’s risk. For example, if we know the distribution $P$ of the data, and know that our algorithm $\mathcal{A}$ is given by $k$-nearest-neighbors (kNN) regression, then we can derive a closed-form expression for the risk $R_{P,n}(\mathcal{A})$ theoretically. Clearly, this is not the setting of interest, since in practice we cannot rely on assumptions about the data distribution—but in fact, even if $P$ is unknown, estimating $R_{P,n}(\mathcal{A})$ for a simple algorithm such as kNN is relatively straightforward. In modern settings, however, a state-of-the-art algorithm $\mathcal{A}$ is typically far more complex, and may be too complicated to analyze theoretically. For this reason, in this work we consider black-box procedures, which analyze the algorithm $\mathcal{A}$ via empirical evaluations, rather than theoretical calculations. The intuition is this: if $\mathcal{A}$ is implemented as a function in some software package, our procedure is allowed to call the function (perhaps as many times as we like, or perhaps with a computational budget), but is not able to examine the implementation of $\mathcal{A}$ (i.e., we cannot read the code that defines the function).
@@
Consider any function ${\widehat{T}}$ that takes as input an algorithm $\mathcal{A}$ and a dataset $\mathcal{D}_{N}\in(\mathcal{X}\times\mathcal{Y})^{N}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})\in\{0,1\}$. We say that ${\widehat{T}}$ is a black-box test if it can be defined as follows: [^2] For this definition, we assume measurability of the map $(\mathcal{D}_{N},\zeta^{(1)},\zeta^{(2)},\dots,\zeta)\mapsto{\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$, which inputs the initial data set $\mathcal{D}_{N}$ and the random seeds used throughout the construction of the test, and returns the output of the test ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})\in\{0,1\}$. for some functions $g^{(1)},g^{(2)},\dots,g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
@@
    $(\mathcal{D}^{(1)},\xi^{(1)})=g^{(1)}(\mathcal{D}_{N},\zeta^{(1)}),$
@@
    $(\mathcal{D}^{(r)},\xi^{(r)})=g^{(r)}\big{(}\mathcal{D}_{N},(\mathcal{D}^{(s)})_{1\leq s<r},({\widehat{f}}^{(s)})_{1\leq s<r},(\zeta^{(s)})_{1\leq s\leq r},(\xi^{(s)})_{1\leq s<r}\big{)},$
@@
    Repeat until some stopping criterion is reached; let ${\widehat{r}}$ denote the (data-dependent) total number of rounds.
@@
    ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=g\big{(}\mathcal{D}_{N},(\mathcal{D}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}^{(r)})_{1\leq r\leq{\widehat{r}}},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big{)}.$
In this definition, at each stage $r$ we are allowed to deploy the training algorithm $\mathcal{A}$ on a data set $\mathcal{D}^{(r)}$ that we have designed adaptively—e.g., by subsampling from the original data set $\mathcal{D}_{N}$, or by generating simulated data given the past information. The random seeds $\zeta^{(1)},\zeta^{(2)},\ldots,\zeta$ allow for randomization at each step of the procedure, if desired—for instance, to allow for randomly subsampling from the original data set $\mathcal{D}_{N}$. (Of course, we might choose to run a deterministic procedure, as a special case.)
An example of a black-box test.
As an example of the type of test that would satisfy Definition 1, we can consider a cross-validation (CV) based estimate of risk. Assuming that $K$ divides $N$ and $n=N(1-1/K)$ for some integer $K\geq 1$, we can construct an estimator as follows:
  • Construct a partition $[N]=S_{1}\cup\dots\cup S_{K}$ with $|S_{k}|=N/K$ for each $k$.
@@
An example that is not black-box.
@@
is possible, since this can be done with only a single call to $\mathcal{A}$; even though this supremum implicitly requires infinitely many evaluations of ${\widehat{f}}$, this is permitted since the functions $g^{(1)},g^{(2)},\dots,g$ in Definition 1 are allowed to depend arbitrarily on the fitted models ${\widehat{f}}^{(1)},{\widehat{f}}^{(2)},\dots$. We make this distinction because, in practice, it is often the case that $\mathcal{A}$ is complex but a typical fitted model ${\widehat{f}}$ is simple—e.g., a sophisticated variable selection procedure that results in a linear regression model.)

### v4

@@ Abstract
Algorithm evaluation and comparison are fundamental questions in machine learning and statistics—how well does an algorithm perform at a given modeling task, and which algorithm performs best? Many methods have been developed to assess algorithm performance, often based around cross-validation type strategies, retraining the algorithm of interest on different subsets of the data and assessing its performance on the held-out data points. Despite the broad use of such procedures, the theoretical properties of these methods are not yet fully understood. In this work, we explore some fundamental limits for answering these questions with limited amounts of data. In particular, we make a distinction between two questions: how good is an algorithm $\mathcal{A}$ at the problem of learning from a training set of size $n$, versus, how good is a particular fitted model produced by running $\mathcal{A}$ on a particular training data set of size $n$? Our main results prove that, for any test that treats the algorithm $\mathcal{A}$ as a “black box” (i.e., we can only study the behavior of $\mathcal{A}$ empirically), there is a fundamental limit on our ability to carry out inference on the performance of $\mathcal{A}$, unless the number of available data points $N$ is many times larger than the evaluation sample size $n$ of interest. On the other hand, evaluating the performance of a particular fitted model can be easy as long as the loss function is bounded and a holdout data set is available—that is, as long as NnN-n is not too small. We also ask whether an assumption of algorithmic stability might be sufficient to circumvent this hardness result. Surprisingly, we find that the same hardness result still holds for the problem of evaluating the performance of $\mathcal{A}$, aside from a high-stability regime where fitted models are essentially nonrandom. Finally, we also establish similar hardness results for the problem of comparing multiple algorithms.
@@ 1 Introduction
Evaluating the performance of a regression algorithm, and comparing the performance of different algorithms, are fundamental questions in machine learning and statistics (Salzberg,, 1997; Dietterich,, 1998; Bouckaert,, 2003; Demsar,, 2006; Garcia and Herrera,, 2008; Hastie et al.,, 2009; Raschka,, 2018). When performing a data analysis, if the properties of the underlying data distribution are unknown, how can we determine which algorithms would model the data well? For instance, for the task of predicting a real-valued response $Y$, which algorithm would have the smallest possible error?
1.0.1 Algorithm evaluation.
@@
[…]
Writing $^{\widehat{f}}$ to denote the fitted model produced by running algorithm $\mathcal{A}$ on a training data set, a related question is to ask about the performance of the trained model $^{\widehat{f}}$:
[…]
@@
In practice, these two questions are often treated as interchangeable: for instance, a paper might state that “Logistic regression performs well on this problem” (which suggests that the question EvaluateAlg is being addressed), but then justify the statement by using a holdout set to evaluate the fitted model $^{\widehat{f}}$ produced by running logistic regression on a particular data set (which is actually addressing the question EvaluateModel). In this work, we argue that these two questions are not equivalent, and indeed, answering EvaluateAlg may be impossible even in settings where EvaluateModel is trivial to answer.
1.0.2 Algorithm comparison.
@@
[…]
@@
[…]
@@ 1.1 Problem formulation: evaluating an algorithm
    with $\mathcal{Y}=\widehat{\mathcal{Y}}$$\mathbb{R}$.
@@
which is computed with respect to a test point sampled as $(X,Y)\sim P$. Now consider a regression algorithm $\mathcal{A}$, which inputs a data set $\{(X_{i},Y_{i})\}_{i\in[n]}$ where $n]:=\{1,\ldots,n\}$ and returns a fitted model ${\widehat{f}}_{n}$. The fitted model ${\widehat{f}}_{n}$ has risk $R_{P}({\widehat{f}}_{n})$—this captures the expected loss for the future performance of ${\widehat{f}}_{n}$, on test data drawn from $P$. Note we can write
@@
where now the expected value is taken with respect to a random draw $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X,Y)\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$—that is, with respect to a random draw of both the training and test data. In particular, the model risk $R_{P}({\widehat{f}}_{n})$ and algorithm risk $R_{P,n}(\mathcal{A})$ are related as
@@
With this notation in place, we can now make our questions, EvaluateAlg and
EvaluateModel, more precise.
[…]
These two questions are different: a low value of $R_{P,n}(\mathcal{A})$ tells us that $\mathcal{A}$ tends to perform well for distribution $P$, while a low value of $R_{P}({\widehat{f}}_{n})$ tells us that this particular fitted model produced by $\mathcal{A}$ performs well. In particular, the latter statement could result from the fact that either $\mathcal{A}$ is good or the particular data set at hand favors $\mathcal{A}$. Which question is the “right” question to ask, will depend on the context, and we do not claim that one of these questions is more important than the other; instead, the aim of our paper is to understand the distinction between these two questions, and to emphasize that the two questions should not be treated interchangeably.
1.1.1 Defining a hypothesis test.
There are many possible goals we might define for addressing EvaluateAlg or EvaluateModel: for example, we might wish to construct a confidence interval for the risk $R_{P,n}(\mathcal{A})$ or $R_{P}({\widehat{f}}_{n})$, or to perform a hypothesis test assessing whether $R_{P,n}(\mathcal{A})$ or $R_{P}({\widehat{f}}_{n})$ lies above or below some prespecified threshold $\tau$,
$\texttt{EvaluateAlg}:H_{0}:\,R_{P,n}(\mathcal{A})\geq\tau\textnormal{ \ versus \ }H_{1}:\,R_{P,n}(\mathcal{A})<\tau,$ (1)
$\texttt{EvaluateModel}:H_{0}:\,R_{P}({\widehat{f}}_{n})\geq\tau\textnormal{ \ versus \ }H_{1}:\,R_{P}({\widehat{f}}_{n})<\tau.$ (2)
We note that it is not hard to see that EvaluateModel is an easy problem as long as the loss function is bounded and we have a large holdout data. This is because after training model ${\widehat{f}}_{n}$, we can use the holdout data $\mathcal{D}_{\textnormal{hold}}$ to estimate $R_{P}({\widehat{f}}_{n})$ via computing $\frac{1}{|\mathcal{D}_{\textnormal{hold}}|}\sum_{(X_{i},Y_{i})\in\mathcal{D}_{\textnormal{hold}}}\ell({\widehat{f}}_{n}(X_{i}),Y_{i})$; if $|\mathcal{D}_{\textnormal{hold}}|$ is large then this estimate is likely to be accurate. However, even in this bounded setting where EvaluateModel is relatively easy, we will see below that EvaluateAlg is fundamentally hard: our results will establish limitations on the ability of any test to answer the question posed by EvaluateAlg, even in the setting of a bounded loss. To establish such a hardness result, it suffices to consider the problem of performing the hypothesis test (1). This is because, if we have instead constructed a confidence interval $\widehat{C}$ with a guarantee of containing the risk $R_{P,n}(\mathcal{A})$ with probability $\geq 1-\alpha$, we can test (1) at Type I error level $\alpha$ by simply checking whether $\widehat{C}\subseteq[0,\tau)$.
@@ 1.2 Overview of contributions
In this paper, for the problem of algorithm evaluation, we will establish fundamental limits on the ability of any black-box test (that is, any test that studies the behavior of $\mathcal{A}$ empirically rather than theoretically) to perform inference on the hypothesis test (1) even when model evaluation is easy to solve. We will also see that this hardness result can be partially alleviated by adding an assumption of algorithmic stability, but only in the high-stability regime. We will also establish analogous results for the problem of algorithm comparison.
Organization. In Section 2, we will present some background on algorithm testing and on algorithmic stability, in addition to giving an overview of some related work in the literature. Section 3 presents our main results establishing the hardness of testing EvaluateAlg, and establishes that an extremely simple and naive test already achieves (nearly) the maximum possible power. Results under an algorithmic stability assumption are presented in Section 4. We then turn to the problem of algorithm comparison, which is addressed in Section 5. Discussion and conclusions are provided in Section 6. All proofs are deferred to the Supplementary Material (Luo and Barber,, 2025).
@@ 2 Background
In this section, we formalize some definitions and give additional background for studying the properties of algorithms. Formally, for data lying in some space $\mathcal{X}\times\mathcal{Y}$, we define an algorithm $\mathcal{A}$ as a map
@@
That is, $\mathcal{A}$ takes as input $\{(X_{i},Y_{i})\}_{i\in[n]}$, a data set of any size, and returns a fitted model $^{\widehat{f}}_{n}:\mathcal{X}\rightarrow\widehat{\mathcal{Y}}$, which maps a feature value $x$ to a fitted value, or prediction, ${\widehat{y}}={\widehat{f}}_{n}(x)$. For instance, for a real-valued response $Y\in\mathcal{Y}=\mathbb{R}$, we might take $\mathcal{A}$ to be the least squares regression algorithm, leading to linear fitted models ${\widehat{f}}_{n}$. In many settings, commonly used algorithms $\mathcal{A}$ will include some form of randomization—for instance, stochastic gradient descent type methods. To accommodate this setting, we will expand the definition of an algorithm to take an additional argument,
@@
The fitted model ${\widehat{f}}_{n}=\mathcal{A}\big(\{(X_{i},Y_{i})\}_{i\in[n]};\xi\big)$ is now obtained by running algorithm $\mathcal{A}$ on data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, with randomization provided by the argument $\xi\in[0,1]$, which acts as a random seed. [^1] Throughout this paper, we implicitly assume measurability for all functions we define, as appropriate. In the case of the algorithm $\mathcal{A}$, this takes the following form: we assume the measurability of the map $\left((x_{1},y_{1}),\ldots,(x_{n},y_{n}),\xi,x\right)\mapsto[\mathcal{A}(\{(x_{i},y_{i})\}_{i\in[n]};\xi)](x)$. Of course, the general definition (3) of a randomized algorithm also accommodates the deterministic (i.e., non-random) setting—we can simply choose an algorithm $\mathcal{A}$ whose output depends only on the data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, and ignores the second argument $\xi$.
@@ 2.1 Black-box tests
To understand the difficulty of the question EvaluateAlg, we need to formalize the limitations on what information is available to us, as the analyst, for performing inference on an algorithm’s risk. For example, if we know the distribution $P$ of the data, and know that our algorithm $\mathcal{A}$ is given by $k$-nearest-neighbors (kNN) regression, then we can derive a closed-form expression for the risk $R_{P,n}(\mathcal{A})$ theoretically. Clearly, this is not the setting of interest, since in practice we cannot rely on assumptions about the data distribution—but in fact, even if $P$ is unknown, estimating $R_{P,n}(\mathcal{A})$ for a simple algorithm such as kNN is relatively straightforward. In modern settings, however, a state-of-the-art algorithm $\mathcal{A}$ is typically far more complex, and may be too complicated to analyze theoretically. For this reason, in this work we consider black-box procedures (Myers et al.,, 2011), which analyze the algorithm $\mathcal{A}$ via empirical evaluations, rather than theoretical calculations. The intuition is this: if $\mathcal{A}$ is implemented as a function in some software package, our procedure is allowed to call the function (perhaps as many times as we like, or perhaps with a computational budget), but is not able to examine the implementation of $\mathcal{A}$ (i.e., we cannot read the code that defines the function).
@@
Let $\mathfrak{D}=\cup_{m\geq 0}(\mathcal{X}\times\mathcal{Y})^{m}$ denotes the space of data sets of any size. Consider any function $^{\widehat{T}}$ that takes as input an algorithm $\mathcal{A}$ and a dataset $\mathcal{D}\in\mathfrak{D}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A},\mathcal{D})\in\{0,1\}$. We say that $^{\widehat{T}}$ is a black-box test if it can be defined as follows: [^2] For this definition, we assume measurability of the map $(\mathcal{D},\zeta^{(1)},\zeta^{(2)},\dots,\zeta)\mapsto{\widehat{T}}(\mathcal{A},\mathcal{D})$, which inputs the initial data set $\mathcal{D}$ and the random seeds used throughout the construction of the test, and returns the output of the test ${\widehat{T}}(\mathcal{A},\mathcal{D})\in\{0,1\}$. for some functions $g^{(1)},g^{(2)},\dots,g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
@@
    $(\mathcal{D}^{(1)},\xi^{(1)})=g^{(1)}(\mathcal{D},\zeta^{(1)}),$
@@
    $(\mathcal{D}^{(r)},\xi^{(r)})=g^{(r)}\big(\mathcal{D},(\mathcal{D}^{(s)})_{1\leq s<r},({\widehat{f}}^{(s)})_{1\leq s<r},(\zeta^{(s)})_{1\leq s\leq r},(\xi^{(s)})_{1\leq s<r}\big),$
@@
    Repeat until some stopping criterion is reached; let $^{\widehat{r}}$ denote the (data-dependent) total number of rounds.
@@
    ${\widehat{T}}(\mathcal{A},\mathcal{D})=g\big(\mathcal{D},(\mathcal{D}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}^{(r)})_{1\leq r\leq{\widehat{r}}},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big).$
In this definition, at each stage $r$ we are allowed to deploy the training algorithm $\mathcal{A}$ on a data set $\mathcal{D}^{(r)}$ that we have designed adaptively—e.g., by subsampling from the original data set $\mathcal{D}$, or by generating simulated data given the past information. The random seeds $\zeta^{(1)},\zeta^{(2)},\ldots,\zeta$ allow for randomization at each step of the procedure, if desired—for instance, to allow for randomly subsampling from the original data set $\mathcal{D}$. (Of course, we might choose to run a deterministic procedure, as a special case.)
2.1.1 An example of a black-box test.
As an example of the type of test that would satisfy Definition 1, we can consider a cross-validation (CV) (Stone,, 1974; Geisser,, 1975) based estimate of risk. Assuming we have $N$ data points $\{(X_{i},Y_{i})\}_{i\in[N]}$, $K$ divides $N$ and n=N(11/K)n=N(1-1/K) for some integer $K\geq 1$, we can construct an estimator as follows:
  • Construct a partition $N]=S_{1}\cup\dots\cup S_{K}$ with $|S_{k}|=N/K$ for each $k$.
@@
2.1.2 An example that is not black-box.
@@
is possible, since this can be done with only a single call to $\mathcal{A}$; even though this supremum implicitly requires infinitely many evaluations of $^{\widehat{f}}$, this is permitted since the functions $g^{(1)},g^{(2)},\dots,g$ in Definition 1 are allowed to depend arbitrarily on the fitted models ${\widehat{f}}^{(1)},{\widehat{f}}^{(2)},\dots$. We make this distinction because, in practice, it is often the case that $\mathcal{A}$ is complex but a typical fitted model $^{\widehat{f}}$ is simple—e.g., a sophisticated variable selection procedure that results in a linear regression model.)